# Relativistic Jets and Winds in Radio-Identified Supermassive Black Hole Binary Candidates

### v1

@@ 2.3 Emission
$I_{\nu,\Omega}(\chi,\psi,\nu)=(1+z_{S})^{-3}d_{A}^{-2}\int j_{\nu_{z},\Omega}e^{-\tau_{\nu}}\cos\theta(\chi,\psi)dr,$ (26)
where $d_{A}$ is the angular diameter distance corresponding to $z_{S}$ and the integration bounds are $r\in[r_{\min},r_{\max}]$. $r_{\min}$ and $r_{\max}$ are the minimum and maximum radii from the black hole which contribute emission at frequency $\nu$. $r_{\min}$ is set by the radius of the photosphere that is optically thick to frequency $\nu$. $r_{\max}$ is more poorly constrained and may relate to where either the collimating wind ceases or synchrotron emission at frequency $\nu$ can no longer be emitted. We return to $r_{\max}$ in Sec. 3.1.1.
@@
$\begin{split}S_{\nu}&=(1+z_{S})^{-3}d_{A}^{-2}\iiint j_{\nu_{z},\Omega}e^{-\tau_{\nu}}dV.\end{split}$ (27)
@@
$\begin{split}\delta I_{\nu}&=(1+z_{S})^{-3}d_{A}^{-2}\int\left(\delta j_{\nu_{z},\Omega}-j_{\nu_{z},\Omega}\delta\tau_{\nu}\right)e^{-\tau_{\nu}}\cos\theta dr,\end{split}$ (28)
@@
$\begin{split}\delta S_{\nu}&=(1+z_{S})^{-3}d_{A}^{-2}\iiint\left(\delta j_{\nu_{z},\Omega}-j_{\nu_{z},\Omega}\delta\tau_{\nu}\right)e^{-\tau_{\nu}}dV,\end{split}$ (29)
@@ 2.4 Model Properties
Figure 2: The local fractional variation in Doppler factor $\delta D/D$ (top), PD $\delta\Pi/\Pi_{0}$ (middle), and EVPA $\delta PA$ in degrees (bottom) for choices of $\beta_{w}$ at a position ($\chi$, $\psi$) within the jet. We fix $\beta_{H}=0.02$, $\Gamma_{j}=10$, and $i=3.8^{\circ}$.
The beaming and polarization variations from a local patch of the jet are given by eqs. 15, 24, and 25. In Fig. 2, we plot $\delta D/D_{0}$, $\delta\Pi/\Pi_{0}$, and $\delta\text{PA}$ for emission from along the jet axis (i.e. $(\chi,\psi)=(0,0)$ so $\hat{r}_{0}=\hat{z}$) at a single $r$. We set $i=3.8^{\circ}$ as inferred for PKS 2131. The ratio $\beta_{H}/\beta_{w}$ controls the variation because the slower the wind, the more rapidly the jet direction changes. Variation in PD is $\delta\Pi/\Pi_{0}\approx-2(\beta_{H}/\beta_{w})\cot i\cos(\Omega t)$ when $\chi\approx 0^{\circ}$. For this very simple model, note that $\Pi_{0}=0$ for $i=0$ (when $\chi=0$). Consequently, at small $i$ (i.e. when the viewer looks directly down the jet), any observed polarization will be induced by the jet tilting. When $\beta_{H}/\beta_{w}\lesssim i$ as shown in Fig. 2, the PD and EVPA oscillate with orbital period $P$ (like the Doppler factor and flux variation). When $\beta_{H}/\beta_{w}\gtrsim i$, the non-perturbative variation will cause the EVPA to swing by $180^{\circ}$ with period $\sim P/2$. We note that this is merely the local polarization behavior from an individual spot on the jet. These represent upper limits on the polarization variation observable when integrating over the entire emitting region. In PKS 2131, ALMA polarization data show possible EVPA rotation by $\sim 20^{\circ}$, while the PD varies from $\sim 4\%$ to $8\%$. Max PD is coincident with minimum flux (Kiehlmann et al., 2025). This is compatible with our model if the contribution of several emission zones smears out the variation.
@@
Integrating eq. 29 to $r_{\max}\gg\lambda_{h}$ can smear out the variation. Consequently, large $\lambda_{h}$ leads to stronger flux variation. The final prediction of this model is that sinusoidal variation will be out of phase at different observing frequencies due to the frequency dependent optical depth (i.e. $\tau\propto\nu^{-\frac{(p+4)}{2}}$). Higher frequencies can thus contribute emission at smaller $r$, causing their light curves to be shifted ahead of lower frequencies. Along with $\beta_{w}$, $r_{\max}$ determines the precise phase shift trend and is likely related to the extent of the wind collimation or particle acceleration inside the jet.
The model we sketch here can be readily generalized to more specific observations of these sources (e.g. polarization and multi-wavelength imaging studies) as well as future-identified sources. Targeted numerical simulations can also enhance our model. Relativistic magnetohydrodynamics (RMHD) simulations can provide more realistic descriptions of the jet geometry, magnetic field structure, and wind dynamics, while kinetic simulations can capture the particle acceleration along the jet.
@@ 3 Interpretation of Observations
Figure 3: PKS 2131-021 light curves at radio, $m$, sub-$m$, and optical wavelengths. (a) Haystack 2.7–31.4 GHz observations (O’Dea et al., 1986) together with the least-squares sine-wave fits to the data at each of the frequencies. (b) OVRO 15 GHz light curve plus ALMA light curves at 91.5, 104, and 345 GHz. (c) Comparison of the $i$, $r$, and $g$-band ZTF optical and OVRO radio light curves of PKS 2131. The curves show the least squares sine wave fits to the corresponding data. (d) The observed phase shifts, relative to the OVRO 15 GHz light curve. A positive phase shift indicates that the light curve is shifted to a later time than the 15 GHz light curve. The curve shows a quadratic polynomial fit to the phase offset. Reproduced from Fig. 6 in Kiehlmann et al. (2025).
@@ 3.1 PKS 2131-021
The $P=1739.8\pm 17.4$ day sinusoidal variation in PKS 2131 has $5.06\sigma$ significance thanks to the multi-wavelength observations across multiple epochs (Kiehlmann et al., 2025). The most well-sampled light curve (OVRO 15 GHz) has sinusoidal flux density variations with amplitude $\sim 0.5$ Jy. There is tantalizing evidence for a higher harmonic with $P/2$ whose amplitude is $\sim 12$ mJy. The fundamental harmonic is significantly identified in 2.7-31.4 GHz Haystack data, 15 GHz OVRO data, 91.5-345 GHz ALMA observations, and ZTF gri data. Fig. 3 shows the light curves in each of the aforementioned bands. Recent ACT light curves complement these datasets in the mm band (Hincks et al., 2025). The periodicity shuts off from 1984-2003 but returns with the same period and the same phase to within the uncertainties (O’Neill et al., 2022). The persistence of the sinusoidal variation is very likely caused by inherent periodicity in the source, consistent with the orbital motion of an SMBHB.
The sinusoidal light curves at low observing frequencies lag behind high observing frequencies. The phase shifts as a function of frequency can be well fitted by a power-law $\delta\phi\propto\nu_{o}^{-n}$, where $n=0.3$ (See bottom right panel in Fig. 3 and discussion in Hincks et al. 2025).
@@ 3.1.1 Model Application
Figure 4: Our model for PKS 2131. (Top) The predicted intensity map at 230 GHz. (Bottom left) The model flux density in a selection of observed bands compared with OVRO and ALMA light curves shifted down by a constant value to match the model. (Bottom Right) Model phase shifts relative to the 15 GHz light compared to the observed phase shifts. Parameters are $\beta_{w}=0.9$, $\beta_{H}=0.02$, $\theta_{j}=0.1$, $\Gamma_{j}=10$, $r_{\min}$ chosen so that $\tau_{\nu}(r_{\min})=20$, $r_{\max}=r_{\min}+60\,(\nu/15\text{ GHz})^{-1}$ ly and $i=3.8^{\circ}$. We assume $L_{e}=10^{46}\text{ erg s}^{-1}\,(r/r_{0})/(1+(r/r_{0}))$ (where $r_{0}=5$ ly) and $L_{m}=10^{46}\text{ erg s}^{-1}\,1/(1+(r/r_{0}))$.
We apply our model described in Sec. 2 to interpret the observations and reproduce the observed sinusoidal variation in PKS 2131. We choose $i=3.8^{\circ}$, consistent with VLBI measurements of the large-scale jet (Homan et al., 2021). For the base jet, we take $L_{e}=10^{46}\text{ erg s}^{-1}\,(r/r_{0})/(1+(r/r_{0}))$ (where $r_{0}=5$ ly), $L_{m}=10^{46}\text{ erg s}^{-1}\,1/(1+(r/r_{0}))$, $\theta_{j}=0.1$, and $\Gamma_{j}=10$. These high jet powers are required to produce $\sim 1$ Jy radio flux at $z_{S}=1.3$, while the scaling is chosen so that $L_{j}=L_{m}+L_{e}$ remains constant. We assume the particle spectrum is initialized with $\gamma_{e,\min}=10$, $\gamma_{e,\max}=300$, and $p=1.2$ and evolves adiabatically along the jet. We return to the spectra later in the discussion. We choose $\beta_{w}=0.9$, $\beta_{H}=0.02$, and $r_{\max}=r_{\min}+60\,(\nu/15\text{ GHz})^{-1}$ ly. Fig. 5 shows the emissivity as a function of radius for the $15$ GHz, $91.5$ GHz, and $345$ GHz emission. In this model, the average radius of the 15 GHz emission is about 20 ly or 10 helical wavelengths. Without the wind, the phase shifts would require the average radius to be 100-500 ly (depending on the jet bulk Lorentz factor), so the emission regions would be separated by factors of several to 10 times more helical wavelengths.
@@
Fig. 4 shows the results. The top panel shows the theoretical 230 GHz intensity map at different phases in the orbit, while the bottom left and right panels show the computed time varying flux density at several observed radio frequencies, and the phase shifts with frequency, respectively. The theoretical curves shown in the bottom left of Fig. 4 are the total flux density predicted by the model. We subtract a constant flux from each real light curve shown so that the average flux of the model matches the observed average flux. We reproduce the observed variation amplitudes while under-predicting the average flux in most bands. We must subtract $\sim 1.1$ Jy from the observed OVRO 15 GHz light curve, $\sim 0.55$ Jy from the observed ALMA 91.5 and 103 GHz light curves, and $\sim 0.3$ Jy from the observed ALMA 345 GHz light curve. These excess fluxes likely come from non-sinusoidally varying large-scale ($\gtrsim 100$ pc) jet components (O’Neill et al., 2022). These non-varying components are observed in the VLBI maps and dominate during the epoch when the periodicity shuts off. At these scales, it is likely that the jet no longer remains a coherent helix, so periodicity is smeared out. Modeling the large-scale components in detail is beyond the scope of this paper.
@@
The imposed $r_{\max}\propto\nu^{-1}$ scaling matches both the sinusoidal amplitudes and phase shifts. Concretely, the $r_{\max}\propto\nu^{-1}$ trend corresponds to a frequency dependent $r$ cutoff in $j_{\nu,\Omega}(r)$ and, consequently, a global cutoff in the non-thermal particle spectrum at maximum Lorentz factor $\gamma_{e,\max}\sim 300$ as assumed. Magnetic reconnection efficiently accelerates particles to energies $\gamma\sim\sigma=B^{2}/4\pi mnc^{2}$ with a very hard spectrum ($p<2$ as we have assumed) (Sironi & Spitkovsky, 2014). The magnetization $\sigma$ is the ratio of magnetic energy to energy in a particular particle species with mass mm and number density $n$. If reconnection is the electron/positron acceleration mechanism in these blazars, this implies $\sigma_{e}\sim 300$. In the emitting regions, the ions likely carry a significant amount of energy ($\sigma_{i}\approx(m_{e}/m_{i})\sigma_{e}\sim 0.3$), which can occur in the boundary layers where there is strong dissipation of the background field. The observed flux may be dominated by boundary layer emission. If further observational and numerical studies point toward this as well, our model can easily be modified so that $L_{e}$ depends on $\chi$.
@@ 3.2 PKS J0805-0111
The $P=1240.7\pm 4.6$ day periodicity in the PKS J0805 OVRO light curve has $3.82\sigma$ significance over the observational epoch from 2008-2022 (de la Parra et al., 2025). The flux density varies by $\sim 0.2$ Jy around an average flux density $0.4$ Jy. Contemporaneous ACT light curves at 95 GHz, 147 GHz, and 225 GHz also show sinusoidal variations. The OVRO light curve lags the higher frequency ACT data just as for PKS 2131 (Hincks et al., 2025), with $\delta\phi\propto\nu_{o}^{-0.7}$. The periodicity ceases at MJD 59041 and has not been observed since. Possible periodic gamma-ray flares are observed 238 days out of phase with the three most recent 15 GHz peaks (de la Parra et al., 2025).

### v2

@@ 2.3 Emission
$I_{\nu,\Omega}(\chi,\psi,\nu)=(1+z_{S})\,d_{L}^{-2}\int j_{\nu_{z},\Omega}e^{-\tau_{\nu}}\cos\theta(\chi,\psi)dr,$ (26)
[…]
@@
$\begin{split}S_{\nu}&=(1+z_{S})\,d_{L}^{-2}\iiint j_{\nu_{z},\Omega}e^{-\tau_{\nu}}dV.\end{split}$ (27)
@@
$\begin{split}\delta I_{\nu}&=(1+z_{S})\,d_{L}^{-2}\int\left(\delta j_{\nu_{z},\Omega}-j_{\nu_{z},\Omega}\delta\tau_{\nu}\right)e^{-\tau_{\nu}}\cos\theta dr,\end{split}$ (28)
@@
$\begin{split}\delta S_{\nu}&=(1+z_{S})\,d_{L}^{-2}\iiint\left(\delta j_{\nu_{z},\Omega}-j_{\nu_{z},\Omega}\delta\tau_{\nu}\right)e^{-\tau_{\nu}}dV,\end{split}$ (29)
@@ 2.4 Model Properties
[…]
The beaming and polarization variations from a local patch of the jet are given by eqs. 15, 24, and 25. In Fig. 2, we plot $\delta D/D_{0}$, $\delta\Pi/\Pi_{0}$, and $\delta\text{PA}$ for emission from along the jet axis (i.e. $(\chi,\psi)=(0,0)$ so $\hat{r}_{0}=\hat{z}$) at some $r$. The precise value of $r$ does not affect the relative phase between $\delta D/D_{0}$, $\delta\Pi/\Pi_{0}$, and $\delta\text{PA}$ and only sets the absolute phase. We set $i=3.8^{\circ}$ as inferred for PKS 2131. The ratio $\beta_{H}/\beta_{w}$ controls the variation because the slower the wind, the more rapidly the jet direction changes. Variation in PD is $\delta\Pi/\Pi_{0}\approx-2(\beta_{H}/\beta_{w})\cot i\cos(2\pi t/P)$ when $\chi\approx 0^{\circ}$. For this very simple model, note that $\Pi_{0}=0$ for $i=0^{\circ}$ (when $\chi=0^{\circ}$), corresponding to no preferred field direction. Consequently, at small $i$ (i.e. when the viewer looks directly down the jet), $\delta\Pi$ contains all the induced polarization and is not perturbative. When $\beta_{H}/\beta_{w}\lesssim i$ as shown in Fig. 2, the PD and EVPA oscillate with orbital period $P$ (like the Doppler factor and flux variation). When $\beta_{H}/\beta_{w}\gtrsim i$, the non-perturbative variation will cause the EVPA to swing by $180^{\circ}$ with period $\sim P/2$. We note that this is merely the local polarization behavior from an individual spot on the jet. These represent upper limits on the polarization variation observable when integrating over the entire emitting region. In PKS 2131, ALMA polarization data show possible EVPA rotation by $\sim 20^{\circ}$, while the PD varies from $\sim 4\%$ to $8\%$. Max PD is coincident with minimum flux (2024Paper2). This is compatible with our model if the contribution of several emission zones smears out the variation.
@@
Integrating eq. 29 to $r_{\max}\gg\lambda_{h}$ can smear out the variation. Consequently, large $\lambda_{h}$ leads to stronger flux variation. The final prediction of this model is that sinusoidal variation will be out of phase at different observing frequencies due to the frequency dependent optical depth (i.e. $\tau_{\nu}\propto\nu^{-\frac{(p+4)}{2}}$). Higher frequencies can thus contribute emission at smaller $r$, causing their light curves to be shifted ahead of lower frequencies. Along with $\beta_{w}$, $r_{\max}$ determines the precise phase shift trend and is likely related to the extent of the wind collimation or particle acceleration inside the jet.
The model we sketch here can be readily generalized to more specific observations of these sources (e.g. polarization and multi-wavelength imaging studies) as well as future-identified sources. Targeted numerical simulations will enhance our model. Relativistic magnetohydrodynamics (RMHD) simulations can provide more realistic descriptions of the jet geometry, magnetic field structure, and wind dynamics, while kinetic simulations can capture the particle acceleration and cooling along the jet.
@@ 3 Interpretation of Observations
Figure 3: PKS 2131-021 light curves at radio, $m$, sub-$m$, and optical wavelengths. (a) Haystack 2.7–31.4 GHz observations (1986AJ.....92.1262O) together with the least-squares sine-wave fits to the data at each of the frequencies. (b) OVRO 15 GHz light curve plus ALMA light curves at 91.5, 104, and 345 GHz. (c) Comparison of the $i$, $r$, and $g$-band ZTF optical and OVRO radio light curves of PKS 2131. The curves show the least squares sine wave fits to the corresponding data. (d) The observed phase shifts, relative to the OVRO 15 GHz light curve. A positive phase shift indicates that the light curve is shifted to a later time than the 15 GHz light curve. The curve shows a quadratic polynomial fit to the phase offset. Reproduced from Fig. 6 in 2024Paper2.
@@ 3.1 PKS 2131-021
The $P=1739.8\pm 17.4$ day sinusoidal variation in PKS 2131 has $5.06\sigma$ significance thanks to the multi-wavelength observations across multiple epochs (2024Paper2). The most well-sampled light curve (OVRO 15 GHz) has sinusoidal flux density variations with amplitude $\sim 0.5$ Jy. There is tantalizing evidence for a higher harmonic with $P/2$ whose amplitude is $\sim 12$ mJy. The fundamental harmonic is significantly identified in 2.7-31.4 GHz Haystack data, 15 GHz OVRO data, 91.5-345 GHz ALMA observations, and ZTF gri data. Fig. 3 shows the light curves in each of the aforementioned bands. Recent ACT light curves complement these datasets in the mm band (2025arXiv250404278H). The periodicity shuts off from 1984-2003 but returns with the same period and the same phase to within the uncertainties (2022ApJ...926L..35O). The persistence of the sinusoidal variation is very likely caused by inherent periodicity in the source, consistent with the orbital motion of an SMBHB.
The sinusoidal light curves at low observing frequencies lag behind high observing frequencies. The phase shifts as a function of frequency can be well fitted by a power-law $\delta\phi\propto\nu_{o}^{-n}$, where $n=0.3$ (See bottom right panel in Fig. 3 and discussion in 2025arXiv250404278H).
@@ 3.1.1 Model Application
Figure 5: Our model for PKS 2131. (Top) The predicted intensity map at 230 GHz. (Bottom left) The model flux density in a selection of observed bands compared with OVRO and ALMA light curves shifted down by a constant value to match the model. (Bottom Right) Model phase shifts relative to the 15 GHz light compared to the observed phase shifts. Parameters are $\beta_{w}=0.9$, $\beta_{H}=0.02$, $\theta_{j}=0.1$, $\Gamma_{j}=10$, $r_{\min}$ chosen so that $\tau_{\nu}(r_{\min})=20$, $r_{\max}=r_{\min}+60\,(\nu(1+z_{S})/15\text{ GHz})^{-1}$ ly and $i=3.8^{\circ}$. We assume $L_{e}=10^{46}\text{ erg s}^{-1}\,(r/r_{0})/(1+(r/r_{0}))$ (where $r_{0}=5$ ly) and $L_{m}=10^{46}\text{ erg s}^{-1}\,1/(1+(r/r_{0}))$.
We apply our model described in Sec. 2 to interpret the observations and reproduce the observed sinusoidal variation in PKS 2131. We choose $i=3.8^{\circ}$, consistent with VLBI measurements of the large-scale jet (2021ApJ...923...67H). For the base jet, we take $L_{e}=10^{46}\text{ erg s}^{-1}\,(r/r_{0})/(1+(r/r_{0}))$ (where $r_{0}=5$ ly), $L_{m}=10^{46}\text{ erg s}^{-1}\,1/(1+(r/r_{0}))$, $\theta_{j}=0.1$, and $\Gamma_{j}=10$. These high jet powers are required to produce $\sim 1$ Jy radio flux at $z_{S}=1.3$, while the scaling is chosen so that $L_{j}=L_{m}+L_{e}$ remains constant. We assume the particle spectrum, whose normalization evolves along the jet, is initialized with $\gamma_{e,\min}=10$, $\gamma_{e,\max}=300$, and $p=1.2$. We return to the spectra later in the discussion. We choose $\beta_{w}=0.9$, $\beta_{H}=0.02$, and $r_{\max}=r_{\min}+60\,(\nu(1+z_{S})/15\text{ GHz})^{-1}$ ly. Fig. 4 shows the emissivity as a function of radius for the $15$ GHz, $91.5$ GHz, and $345$ GHz emission. In this model, the average radius of the 15 GHz emission is about 20 ly or 10 helical wavelengths. Without the wind, the phase shifts would require the average radius to be 100-500 ly (depending on the jet bulk Lorentz factor), so the emission regions would be separated by factors of several to 10 times more helical wavelengths.
Fig. 5 shows the results. The top panel shows the theoretical 230 GHz intensity map at different phases in the orbit, while the bottom left and right panels show the computed time varying flux density at several observed radio frequencies, and the phase shifts with frequency, respectively. The theoretical curves shown in the bottom left of Fig. 5 are the total flux density predicted by the model. We subtract a constant flux from each real light curve shown so that the average flux of the model matches the observed average flux. We reproduce the observed variation amplitudes while under-predicting the average flux in most bands. We must subtract $\sim 1.1$ Jy from the observed OVRO 15 GHz light curve, $\sim 0.55$ Jy from the observed ALMA 91.5 and 103 GHz light curves, and $\sim 0.3$ Jy from the observed ALMA 345 GHz light curve. These excess fluxes likely come from non-sinusoidally varying large-scale ($\gtrsim 100$ pc) jet components (2022ApJ...926L..35O). These non-varying components are observed in the VLBI maps and dominate during the epoch when the periodicity shuts off. At these scales, it is likely that the jet no longer remains a coherent helix, so periodicity is smeared out. Modeling the large-scale components in detail is beyond the scope of this paper.
@@
The imposed $r_{\max}\propto\nu^{-1}$ scaling matches both the sinusoidal amplitudes and phase shifts. Concretely, the $r_{\max}\propto\nu^{-1}$ trend corresponds to a frequency dependent $r$ cutoff in $j_{\nu,\Omega}(r)$ and, consequently, a global cutoff in the non-thermal particle spectrum at maximum Lorentz factor $\gamma_{e,\max}\sim 300$ as assumed. Up to now, we have not made any assumptions about the acceleration mechanism needed to sustain our postulated non-thermal spectrum. Magnetic reconnection efficiently accelerates particles to energies $\gamma\sim\sigma=B^{2}/4\pi mnc^{2}$ with a very hard spectrum ($p<2$ as we have assumed) (2014ApJ...783L..21S). The magnetization $\sigma$ is the ratio of magnetic energy to energy in a particular particle species with mass mm and number density $n$. If reconnection is the electron/positron acceleration mechanism in these blazars, this implies $\sigma_{e}\sim 300$ in the acceleration zone. In terms of the electromagnetic and radiating particle powers, we have $\sigma\approx L_{m}/L_{e}=(r/r_{0})^{-1}$, which would suggest the dissipation begins at radii $r\sim r_{0}/\sigma_{e}\approx 0.02$ ly for the parameters we have assumed. At this small radius, the particle population would likely be strongly cooled, so the acceleration zone would need to span a range of radii to sustain the particle population. If the magnetic energy is principally dissipated into ions, then acceleration via reconnection can continuously occur at larger radii. In the emitting regions, the ions, which we have not otherwise accounted for, could carry a significant amount of energy, so $\sigma_{e}$ may remain high even as $\sigma_{i}$ shrinks ($\sigma_{i}\approx(m_{e}/m_{i})\sigma_{e}\sim 0.3$). Strong dissipation of the background field and ion mixing can occur near the wind-jet boundary layers, so the observed flux may be dominated by boundary layer emission. If further observational and numerical studies point toward this as well, our model can easily be modified so that $L_{e}$ depends on $\chi$. Edge-brightened emission is likely to come from reconnection-accelerated particles, while emission from the center of the jets is more likely to come from shock-accelerated particles. More work will be required to understand particle acceleration in the jets.
@@ 3.2 PKS J0805-0111
The $P=1240.7\pm 4.6$ day periodicity in the PKS J0805 OVRO light curve has $3.82\sigma$ significance over the observational epoch from 2008-2022 (2025ApJ...987..191D). The flux density varies by $\sim 0.2$ Jy around an average flux density $0.4$ Jy. Contemporaneous ACT light curves at 95 GHz, 147 GHz, and 225 GHz also show sinusoidal variations. The OVRO light curve lags the higher frequency ACT data just as for PKS 2131 (2025arXiv250404278H), with $\delta\phi\propto\nu_{o}^{-0.7}$. The periodicity ceases at MJD 59041 and has not been observed since. Possible periodic gamma-ray flares are observed 238 days out of phase with the three most recent 15 GHz peaks (2025ApJ...987..191D).